please help, I don't know if my answer is correct or not , if not which one? thank youThe controller of Hendershot Corporation estimates the amount of materials handling overhead cost that should be allocated to
Answer #1
material handling overhead rate = total material handling cost/total expected material move
material handling overhead rate = 16562.20/3150 = 5.2578
Total material handling cost allocated to modular homes = 1500*5.2578
Total material handling cost allocated to modular homes = 7887
Option " $7887" Is Correct

Please give a thumbs up if it is helpful & let me know if any doubt